OpenStudy (anonymous):

the equation sec^2x-1=tan^2 x is an identity? true or false need help

OpenStudy (xapproachesinfinity):

try some values if it doesn't hold then it is not an identity otherwise it is an identity

OpenStudy (anonymous):

thing it is identity

OpenStudy (anonymous):

think it is

OpenStudy (thomas5267):

Start with:\[ \cos^2(2\theta)+\sin^2(2\theta)=1\\ \frac{\cos^2(2\theta)}{\cos^2(2\theta)}+\frac{\sin^2(2\theta)}{\cos^2(2\theta)}=\frac{1}{\cos^2(2\theta)}\\ 1+\tan^2(2 \theta)=\sec^2(\theta)\\ \sec^2(2\theta)-1=\tan^2(\theta) \]

OpenStudy (anonymous):

True.